OpenStudy (safia21):

In a right triangle, if one acute angle has measure 6x-10 and the other has measure 2x- 4 , find the larger of these two angles.

OpenStudy (asadkarim7):

the equation would be (6x-10) + (2x-4) +90 = 180 solve for x and then put the value of x in both angles to find which is bigger

OpenStudy (safia21):

how do you solve for x

OpenStudy (asadkarim7):

eliminate the braces in (above posted ) equation and on LHS keep the coefficient value

OpenStudy (anonymous):

You lift the parenthesizes: 8x-14+90=180 , 8x+76=180 Then you isolate the numbers and the x'es: 8x=104 Then you divide by 8 to get x x=13

OpenStudy (asadkarim7):

like 6x -10 +2x -4 = 90 6x +2x = 104 8x = 104 x = 13

OpenStudy (safia21):

thanks can you guys help me with one more problem

OpenStudy (asadkarim7):

sure

OpenStudy (asadkarim7):

link me

OpenStudy (safia21):

∆ABC if BC is one inch longer AB, AC is 10 inches shorter than the sum of BC and and AB and the perimeter of ∆ABC is 72 inches find the length of BC

OpenStudy (asadkarim7):

oh k first assume the length of AB = x Bc would be = x+1 Ac would be = ( x + x+1) -10 = (2x +1) -10 = 2x -9 x + x+1 + 2x-9 = 72 2x +1 + 2x -9 =72 4x -8 =72 4x =80 x = 20 so BC would be x + 1 x= 20 so BC= 21